OpenStudy (anonymous):

A rental truck costs $50 plus $.50 per mile. Ues intersection of graphs method to find out number of miles the truck will drive to cost $80

OpenStudy (anonymous):

it drove 60 miles, unfortunately I can't draw a graph on here.

OpenStudy (anonymous):

But your graphs will be 50 + .5x, and x=80

OpenStudy (anonymous):

Cost= 50+0.5x So you set to the cost=80 80=50+0.5x 30=0.5x 60=x I'm not sure what the intersection of graphs method is....

OpenStudy (anonymous):

Thank you, also how would I solve that numerically with a table of values?